### 03 Mar 2024 v1

# Ever-Evolving Memory by Blending and Refining the Past

Abstract
For a human-like chatbot, constructing a long-term memory is crucial. A naive approach for making a memory could be simply listing the summarized dialogue. However, this can lead to problems when the speaker’s status change over time and contradictory information gets accumulated. It is important that the memory stays organized to lower the confusion for the response generator. In this paper, we propose a novel memory scheme for long-term conversation, CREEM. Unlike existing approaches that construct memory based solely on current sessions, our proposed model blending past memories during memory formation. Additionally, we introduce refining process to handle redundant or outdated information. This innovative approach seeks for overall improvement and coherence of chatbot responses by ensuring a more informed and dynamically evolving long-term memory.
Ever-Evolving Memory by Blending and Refining the Past
1 Introduction
Humans undergo continuous changes in emotions and states over time. For instance, one who was into classical music might turn into a passionate fan of heavy metal, or one who was suffering from severe illness could now be healthy enough to jog every morning. In order to create a chatbot that is sensitive to these changes, it is crucial that the changes are immediately reflected in its long-term memory Adiwardana et al. (2020). In this paper, we introduce a memory model that seamlessly blends and refines past memories. We define two problems to make an advanced long-term memory that resolves the issues of the existing memory frameworks:
Figure 1: Motivating examples where our memory (Blend + Refine) is needed compared to not refining (Case 1) and removing all contradictories (Case 2).
Figure 2: Overall memory update process of CREEM.
[P1] Only the current session is considered when creating new memory.
Long-term memory can be constructed by simply storing the summarized dialogue or persona Jang et al. (2023); Xu et al. (2021). This approach, focusing solely on the current session, results in one-dimensional information, leading to unorganized and disconnected memories. Consequently, the integration of past memories becomes overly dependent on the response-generating language model.
[P2] Memory contains outdated or contradictory information.
Prior works have explored expanding memory beyond direct dialogue information using commonsense knowledge or high-level insights from recent memory Majumder et al. (2020); Kim et al. (2023); Park et al. (2023). However, these methods risk introducing excessive redundancy or contradictions. Figure 1 illustrates Case 1, where outdated memory might be prioritized, neglecting the current context. Additionally, simply removing contradictions Bae et al. (2022) during the refining process may result in missing information, leading to awkward conversations, as depicted in Case 2 of Figure 1.
To address these issues, we propose a novel memory framework Contextualized Refinement based Ever-Evolving Memory (CREEM). Our goal is to make an ever-evolving memory that effectively integrates long-term memory, seamlessly connecting past and new information. CREEM continuously refines its memory every time there is a new information to remember. The main contributions of our work are as follows:
  • We introduce CREEM, a novel memory constructing framework to tackle the issues of fragmented and contradictory memories.
  • We establish new criteria for what constitutes a good memory and evaluate based on them.
  • We demonstrate how the memory quality aligns with response generating performance.
2 Proposed Approach: CREEM
2.1 Tasks
Task 1. Blending Past Memory (P1)
We should take past memories into consideration while making a new memory to be added. Prior works that construct memory with the information extracted from a single session can be expressed as equation 1 where $\mathrm{m}_{s}$ and $\mathbf{D}^{s}$ are new memories and dialogue at current time $s$ respectively, and $f$ is the memory expanding function. On the contrary, our method utilizes past memories while constructing the memory itself as equation 2 where $\mathbf{M}_{[0,s-1]}$ is the set of past memories accumulated.
$\mathrm{m}_{s}=f(\mathbf{D}^{s})$ (1)
$\mathrm{m}_{s}=f(\mathbf{D}^{s},\mathbf{M}_{[0,s-1]})$ (2)
Task 2. Refining Memory (P2)
When adding new memory, past memory should be refined to match the overall consistency. We refine the existing memory by judging whether they are redundant or updated.
In this section, we will explain the details of each step. The overall method flow is shown in Figure 2 and Algorithm 1.
2.2 Contextual Search
The initial step is searching past memories helpful in forming new ones. To search for contextually relevant memories, we formulate a query using ChatGPT-3.5-turbo  OpenAI (2023). The generated question serves to extract top-k information not only from past memories but also the ongoing dialogue, ensuring recent utterances are considered in the memory retrieval process.
Evolutionary Fundamental
Integration Consistency Sophistication Conciseness Relevance Concreteness Diversity
CC Dataset
$\text{SumMem}_{\text{MSC}}$ 3.16 4.11 2.59 3.72 3.85 3.10 3.24
$\textsc{ReBot}_{\text{CC}}$ 4.01 4.50 3.57 3.90 4.51 4.01 4.00
CREEM 4.05 4.64 3.86 3.95 4.61 4.00 4.03
MSC Dataset
$\text{SumMem}_{\text{MSC}}$ 3.44 3.95 3.02 3.25 4.09 3.82 3.72
$\textsc{ReBot}_{\text{CC}}$ 3.86 4.11 3.52 3.36 4.23 3.97 4.00
CREEM 4.02 4.19 3.88 3.38 4.40 4.01 4.07
Table 1: G-eval of memory based on 7 criteria with CC and MSC dataset.
2.3 Blend and Refine
The Blend process uses the top-k information previously retrieved. Given those information, $n$ memories are newly generated through integration and high-level inference.
Simultaneously, the Refine step evaluates the relationship between new and past memories. It categorizes the inference into three options: New if it contains entirely novel information, Redundant if it overlaps with past memories, and Updated if past information has changed. In the case of Redundant or Updated, LLM identifies which past memories to remove. After deletion, the new insights are added, and the model generates a response based on the current dialogue and the added insights.
1: Definition: $GQ=GenerateQuery$,
2:$R=Retrieval$,   $BR=Blend\&Refine$
3: $GR=GenerateResponse$
4: Input: M, D , s , t
5: $query\leftarrow GQ\left(\mathbf{D}^{s}_{[1,t-1]}\right)$
[…]
8:$Response\leftarrow GR\left(\mathbf{M}_{t}^{s},\mathbf{D}^{s}_{[1,t-1]}\right)$
Algorithm 1 Memory Updating Process
CREEM $\textsc{ReBot}_{\text{CC}}$ $\text{SumMem}_{\text{MSC}}$
What present + past present present
When per turn per session per turn
Who each speaker no distinction each speaker
Model ChatGPT (two-shot) distilled T5-base supervised encoder-decoder
Fine-tuned $\times$ $\checkmark$ $\checkmark$
Table 2: Comparison of memory constructing methods.
3 Experiments
3.1 Datasets
We evaluate on two long-term chat datasets. Conversation Chronicles (CCJang et al. (2023) is a dataset generated by ChatGPT given a relationship, time interval and the storyline employ the narratives from SODA Kim et al. (2022). Multi-Session Chat (MSCXu et al. (2021) is a dataset created by human annotators through expanding dialogues from Persona-Chat Zhang et al. (2018), given a set of persona of previous sessions..
We focus on using these datasets mainly as a source of rich topics and conversation settings for our experiment, rather than absolute benchmarks. In order to validate the effect of a long-term memory, we experiment for the final sessions which has accumulated sufficient memory. For testing, we use 5 utterances per turn across 50 episodes from each datasets, yielding approximately 250 test data each.
3.2 Models and Settings
We compare our memory approach with two other methods introduced with the datasets. For the CC dataset, $\textsc{ReBoT}_{\text{CC}}$ summarizes each session’s dialogue with a distilled T5-base model Raffel et al. (2020), using these summaries as long-term memory for both speakers. In the case of the MSC dataset, they employ its summarization model, which we will call as $\text{SumMem}_{\text{MSC}}$, to extract persona for each speaker’s utterance. Table 2 provides a detailed comparison.
We set both baseline methods to receive the entire memory as input for response generation, following the original $\textsc{ReBoT}_{\text{CC}}$ approach. For CREEM, Contriever Izacard et al. (2021) is used as the retriever to extract 5 information pieces, and the Blend process generates 2 new memories. For fairness, we use ChatGPT as the response generation model for all methods.
3.3 Results and Discussion
3.3.1 Memory Evaluation
CREEM makes high quality memory.
First of all, we should define what is a good memory? A good memory should consistently offer relevant information aligned with past conversations. Similar to human memory, it should evolve, keeping track of changes and resolving conflicts between old and updated information.
We establish 4 criteria for evaluating the qualities as a next-generation memory, under the evolutionary aspect: accurate connection and reflection of changes across sessions in details (Integration); absence of contradictory or conflicting information (Consistency); advancement beyond one-dimensional, direct information (Sophistication); well-organized, avoiding unnecessary repetition or extraneous details (Conciseness).
We set additional 3 criteria to assess the fundamental aspects of a memory: close relation to the speaker (Relevance); specific and precise information (Concreteness); variety of topics, perspectives, or insights (Diversity).
We conduct G-Eval Liu et al. (2023) customized to evaluate the memory given each of the criteria, instruction steps, and the conversation of all sessions.
Table 1 displays scores for three models across datasets. CREEM outperforms in all four evolutionary aspects, indicating its superior ability to integrate, maintain consistency, achieve sophistication, and ensure conciseness. The Blend process appears to play a crucial role in integrating information across sessions, creating sophisticated memories by expanding the memory. Additionally, the Refine process contributes to maintaining a consistent and well-organized memory by continually removing redundant or outdated information. Even in fundamental aspects, CREEM excelled in all criteria except for a slight difference in Concreteness in the CC dataset.
CREEM vs. $\textsc{ReBoT}_{\text{CC}}$ $\text{SumMem}_{\text{MSC}}$
CC Dataset
   Consistency 61.29 % 74.52 %
   memorability 48.06 % 70.97 %
MSC Dataset
   Consistency 63.85 % 67.57 %
   memorability 61.15 % 61.15 %
Table 3: Comparison of generated response quality. We report CREEM’s winning rate.
CREEM
Original vs. ()(-)( - ) refine ()(-)( - ) pre-rg Blend
CC Dataset
   Consistency 64.41 % 53.55 %
   memorability 59.66 % 47.74 %
MSC Dataset
   Consistency 64.32 % 63.77 %
   memorability 63.40 % 58.70 %
Table 4: Results of ablation study. We report CREEM’s winning rate.
3.3.2 Response Generation Evaluation
CREEM’s memory improves response generation.
We conduct a pair-wise evaluation with ChatGPT based on 2 criteria that can most effectively show the role of the memory: maintains consistency without contradictions from past sessions (Consistency); effectively recalls past events correctly by retaining information from previous sessions (Memorability).
Table 3 shows the winning rate of CREEM. With CREEM’s memory, LLM creates better response in terms of consistency. $\textsc{ReBoT}_{\text{CC}}$ seems to slightly excel in memorability within it’s own dataset. This is possible considering that the CC dataset itself is made with ChatGPT and $\textsc{ReBoT}_{\text{CC}}$ uses fine-tuned memory model.
3.3.3 Ablation Study
Refinement has a crucial effect both in memory and response quality.
Table 4 shows the performance difference between CREEM and its variants excluding refine processes. We evaluate their consistency and memorability against CREEM origin responses in the same way as above. Overall, it was observed that without refining, the responses that are less consistent and has worse memory.
Pre-response blending helps generate better response.
Table 4 shows the performance difference between CREEM and its variants excluding blending right before generating reponse. This setting only gives memories accumulated in the past session, while CREEM’s original method also provides new memories created in the ongoing session until its responding turn. Overall, CREEM consistently produces more consistent and memory-enhanced responses. However, with CC data, CREEM without pre-RG blending is slightly favored in terms of memorability, possibly due to limited retrieval or blend creations. Increasing these hyperparameters is expected to address this issue.
4 Conclusion
We created an ever-evolving memory scheme to incorporate both current and past memories, actively refining outdated or contradictory information. Despite using a maximum 2-shot LLM prompting approach, in contrast to other fine-tuned methods, CREEM showed superior memory alignment and response generation. Through CREEM, we aim to create a more human-like memory system, fostering further evolution of chat-bot memory systems.
5 Limitation
Our study has the following limitations: (1) Apart from the scheme of CREEM, the results can be influenced by the performance of the retriever. As future work, we plan to develop a retriever model specialized in memory and conversation data; (2) During the Blend & Refine process, CREEM operates within the top-k retrieved memories. However, it may overlook removing insights or updated relationships when the disappearing past memories are less than or equal to k. Currently, our method doesn’t handle unretrieved memories, but incorporating graph relationships among memories in the future could enhance memory retrieval and management; (3) There is not enough publicly open high-quality long-term datasets. We expect CREEM to have an advantage over other models as the conversations become very long through the blend and refine process. We plan to experiment by developing a longer dataset for better memory evaluation.
Appendix A Method Prompts
A.1 Query Prompt
We use ChatGPT to create an appropriate query that could be generated for all conversations from the current session. We use a two-shot prompt for this task, but we don’t include all examples due to the length. The prompt for this task is as follows:
You will act as a model that generates a query relevant to the given dialogue.
Given the dialogue, what is the most salient high-level question we can answer about speaker, focusing mainly on the last utterance?
Dialogue:{text}
Question about {speaker}:
A.2 Blend for first session Prompt
We use ChatGPT to generate two insights for the first session. For this task, a one-shot prompt is used, and examples are omitted for length. The prompt is as follows:
You need to act as a memory expansion model.
Given a set of statements, you should generate 5 high-level insights about {speaker}.
<Statements>:
{text}
<Results about {speaker}>:
A.3 Blend and Refine Prompt
We use ChatGPT to generate two insights related to insights and utterances that have been retrieved for queries in a second or later session. For this task, a two-shot prompt is used, and examples are omitted for length. The prompt is as follows:
You need to act as a memory expansion model following two steps.
First, given a set of statements, you should generate high-level insights about {speaker}.
Next, with the high-level insights, you should refine the given statements.
Determine whether the new insight is [New], [Redundant], or [Updated] from the given statements. If it is [Redundant] or [Updated], indicate which statements to remove. If it is [New], there are no statements to remove.
Continue these steps twice to make two meaningful insights.
Here are some examples that can be helpful.
<Statements>:
{text}
<Results about {speaker}>:
A.4 Respond Prompt
After the Refine stage, ChatGPT generate an appropriate response to the previous utterance based on the newly established memory. Three guidelines are given to create an appropriate response. The prompt is as follows:
You will act as a response generator. Considering the given [Memory], continue the given conversation and create an appropriate response for the last utterance.
Consider the following factors when generating:
- Do not make the response too long.
- Do not contain any of the speaker’s name explicitly.
- Reflect the [Memory] in the generated response.
[Memory of {speaker}]:
{memory_stream}:
[Conversation]:
{current_dialogue}: [Response] {speaker}:
Appendix B Eval Prompts
B.1 Response Evaluation Prompt
We use ChatGPT to pairwise compare the responses generated by CREEM with the responses generated by $\text{SumMem}_{\text{MSC}}$ and $\textsc{ReBoT}_{\text{CC}}$. In conducting this work, two criteria are used: consistency and memorability. An example of the format is given and the criteria were specified. The criteria according to the evaluation criteria are as follows.
  • Consistency: The response should maintain consistency without contradictions from past sessions. The response aligns seamlessly with what has been discussed in previous sessions, following the individual’s characteristic and current status.
  • Memorability: The response should effectively recalls past events correctly by retaining information from previous sessions. The response actively uses memory from the past dialogue to continue the conversation.
Considering that the results are slightly different each time ChatGPT generates, a total of 19 results were generated at once, and the one selected with more was taken as the result. Among the two prompts, the following is one that evaluates consistency. Format examples have been omitted due to length:
You will be given a conversation between two individuals. You will then be given two response options for the next turn in the conversation.
Your task is to choose the better response based on one metric with a brief explanation.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
Evaluation Criteria:
Consistency - The response should maintain consistency without contradictions from past sessions. The response aligns seamlessly with what has been discussed in previous sessions, following the individual’s characteristic and current status.
Now choose the response that has better consistency given the past and current session dialogue.
Past Sessions:
{past_dialogue}
Current Session:
{current_dialogue}
Response Options:
{response}
Better Reponse:
B.2 Memory Evaluation Prompt
We compare which of the memories created by $\text{SumMem}_{\text{MSC}}$, $\textsc{ReBoT}_{\text{CC}}$, and CREEM had better performance through ChatGPT using the CC dataset and the MSC dataset. There are a total of 7 criteria used in this evaluation, and each prompt contains different evaluation steps for each criteria so that we can evaluate which memory better reflects the conversation by looking at the total dialogue and the memory created for the dialogue. Considering that the results are slightly different each time ChatGPT generates, a total of 20 results are generated at a time and the average is taken as the result. The evaluation step is omitted due to length. The seven criteria are as follows:
  • Integration (1-5): This criterion evaluates whether the memory effectively connects information across sessions and accurately reflects changes in details about {speaker} over time. The goal is to prioritize memories that present a cohesive organization of content across sessions.
  • Consistency (1-5): This criterion evaluates whether there are any contradictory or conflicting pieces of information about {speaker} within the memory. The goal is to ensure that the information presented about {speaker} remains coherent and free from contradictions throughout the conversation.
  • Sophistication (1-5): This criterion assesses whether the content of the memory is advanced and goes beyond one-dimensional, direct information about {speaker} from the conversation.
  • Conciseness (1-5): This criterion evaluates whether the memory is well-organized, avoiding unnecessary repetition or extraneous information.
  • Relevance (1-5): This criterion evaluates whether the memory is closely related to the {speaker} of the conversation. It ensures that the content contributes meaningfully to the ongoing conversation.
  • Concreteness (1-5): This criterion evaluates whether the memory contains substantial and specific information about {speaker}. It assesses the level of detail and precision in the representation of {speaker}.
  • Diversity (1-5): This criterion assesses whether the content of the memory has diverse information about {speaker}.
Appendix C Examples of Memory construction and Response generation
Table 6 shows an example of building a memory and generating an answer using the methods of $\text{SumMem}_{\text{MSC}}$, $\textsc{ReBoT}_{\text{CC}}$, and CREEM for one episode from a random MSC dataset.
Redundant New Updated
MSC dataset 24.61 % 55.81 % 19.59 %
CC dataset 30.35 % 53.06 % 16.59 %
Table 5: Ratio of Redundant, New, Updated during constructing memory by CREEM
Dialogue Context (msc-episode-40-s4-t5)
Speaker A: Well, my hair has been dyed!
Speaker B: How did it come out? I bet your grandkids were thrilled.lol
is her passion, she was an absolute natural with keeping my nerves at bay.
Speaker B: At least it’s less permanent than a drastic hair cut would be.
You’ve got extra flavor now. Did you stick with just a streak or did you go bigger?
Speaker A: We decided to do a streak on each side, for balance. And they’re
not just out in the open but under a layer of hair, so it pops out when I wear
a ponytail. My granddaughter suggested that, how creative!
$\text{SumMem}_{\text{MSC}}$
    Memory I am excited for Easter Weekend. I have blue hair. I have a new Easter dress.
I am 29 years old. I am going on spring break. My mom just gave me a new car. I drive.
I dyed my hair with henna. I love grey hair. I want to dye my hair grey. My dress is turquoise.
My hair is blue. I just got my first new car. I’ve been driving pre owned cars before.
I have red hair. I have a friend. I have a VW Tiguan. I fill up my car once a month.
I don’t drive a lot. I work as a librarian at the beach.
    Response That sounds like a fun and unique way to add some color to your hair!
It’s great that your granddaughter was able to offer a creative suggestion. I’m sure it looks amazing
when you wear a ponytail.
$\textsc{ReBoT}_{\text{CC}}$
    Memory A retired librarian plans to leave for spring break due to her mother’s new car
and quiet nature. She enjoys nature, but is retired due to retiring.
A woman dyes her hair red with henna, which is good for red hair. She also suggests
doing a semi-permanent dye and having fun with her grandkids if allowed .
Turquoise is her favorite color, but her new dress is turquoise.
A grandkid dyed her hair turquoise, which she thinks will look great on her youngest
grandson’s favorite color. They plan to touch up their hair for the summer months due to
the sun and pool. They also have recommendations for a day visit at the beach."
    Response That sounds like a fun and unique way to add some color to your hair! I’m glad your
granddaughter had such a creative idea. It must be exciting to have a little pop of
color that shows when you wear a ponytail."
CREEM
    Memory Speaker B is open to experimenting with different hair colors and is specifically interested
in dyeing their hair grey in the future. Speaker B highly values the opinions and
expertise of their grandkids when it comes to hair dyeing decisions.
    Response Your grandkids really have a good eye for hair styling! I remember you mentioning that
you were interested in dyeing your hair grey in the future. Maybe you can consult
them for some ideas when the time comes.
Table 6: Comparison of memory and response generation between CREEM, $\text{SumMem}_{\text{MSC}}$, $\textsc{ReBoT}_{\text{CC}}$ for MSC dataset
Appendix D Analysis of Refine
Table 5 shows the extent to which Redundant, New, and Updated processes occur when creating memory with CREEM through 50 MSC datasets and 50 CC datasets. When using the MSC dataset, the rate of Redundant and Updated occurrences is about 44.2 %, and when using the CC dataset, it is about 46.94 %. Thus, we confirm that the process of checking and refining whether there are redundant or updated memories among past memories was performed properly.
Appendix E Implementation details
E.1 Contriever
In our implementation into CREEM, We use Contriever  Izacard et al. (2021) to retrieve relevant past insights or utterances for queries created after the previous utterance. Contriever, trained through unsupervised contrastive learning, serves as a dense information retriever. Remarkably, even in the absence of explicit supervision, it exhibits outstanding proficiency in information retrieval tasks.
E.2 Large Language Model
In this study, we use ChatGPT for implementing Query Generation, Blend and Refine and Response Generation. ChatGPT, a Large Language Model (LLM) with 175 billion parameters, is built upon Instruct-GPT Ouyang et al. (2022). It is specifically trained to adhere to user instructions and provide requested information in a conversational style. To interact with the OpenAI API, we utilize LangChain.
Appendix F License and Terms
For our implementation and evaluation, we use Huggingface library [^1] https://huggingface.co/ and vLLM library. Both libraries are licensed under Apache License, Version 2.0. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.